OpenStudy (anonymous):

Find the critical points of f(x)=(2^x)(sin(x) . I can only find -.96, but my online homework is telling me there are two more.

OpenStudy (anonymous):

is this \[f(x)=2^x\sin(x)\]?

OpenStudy (anonymous):

derivative is \[2^x(\cos(x)+\ln(2)\sin(x))\] \[2^x>0\] so you have to solve \[\cos(x)+\ln(2)\sin(x)=0\] good luck, but i think there might be several